OpenStudy (anonymous):

log (base 1/3) (27)+log (base 1/3) (1/9) express as a single logarithm, simplify if possible.

OpenStudy (ybarrap):

log(27) + log(1/9), where log is base 1/3 = log (3 * 9) + log (1) - log (9) = log(3) + log(9) + 0 - log(9) = log(3) = -1, since (1/3)^(-1) = 3. Ans. -1.